*Please show excel function*

What is the probability that a student will be over 60 inches tall? (assume that the data are normally distributed with the population standard deviation =4.25)

Height in inches

Student 1 64
2 63
3 63
4 63
5 63
6 65
7 66
8 66
9 67
10 68
11 68
12 66
13 68
14 78
15 76
16 75
17 75
18 73
19 73
20 72
21 70
22 70
23 70
24 70
25 69
26 69
Answer #1
Values ( X ) \Sigma (X_{i} - \bar{X})^{2}
64 23.4857
63 34.1781
63 34.1781
63 34.1781
63 34.1781
65 14.7933
66 8.1009
66 8.1009
67 3.4085
68 0.7161
68 0.7161
66 8.1009
68 0.7161
78 83.7921
76 51.1769
75 37.8693
75 37.8693
73 17.2541
73 17.2541
72 9.9465
70 1.3313
70 1.3313
70 1.3313
70 1.3313
69 0.0237
69 0.0237
Total 1790 465.3858

Mean \bar{X} = \Sigma X_{i} / n

\bar{X} = 1790 / 26 = 68.8462

X \sim N ( \mu = 68.8462 , \sigma = 4.25 )

P ( X > 60 ) = 1 - P ( X < 60 )

Standardizing the value

Z = ( X - \mu ) / \sigma

Z = ( 60 - 68.8462 ) / 4.25

Z = -2.08

P ( ( X - \mu ) / \sigma ) > ( 60 - 68.8462 ) / 4.25 )

P ( Z > -2.08 )

P ( X > 60 ) = 1 - P ( Z < -2.08 )

P ( X > 60 ) = 1 - 0.0188

P ( X > 60 ) = 0.9812

Excel function   1-NORMSDIST( Z = - 2.08 ).